# Don’t You (Project Around Discs)? Neural Network Surrogate and Projected Gradient Descent for Calibrating an Intervertebral Disc Finite Element Model

### v1

@@ Abstract
Accurate calibration of finite element (FE) models of human intervertebral discs (IVDs) is essential for their reliability and application in diagnosing and planning treatments for spinal conditions. Traditional calibration methods are computationally intensive, requiring iterative, derivative-free optimization algorithms that often take hours or days to converge.
This study addresses these challenges by introducing a novel, efficient, and effective calibration method for an L4-L5 IVD FE model using a neural network (NN) surrogate. The NN surrogate predicts simulation outcomes with high accuracy, outperforming other machine learning models, and significantly reduces the computational cost associated with traditional FE simulations. Next, a Projected Gradient Descent (PGD) approach guided by gradients of the NN surrogate is proposed to efficiently calibrate FE models. Our method explicitly enforces feasibility with a projection step, thus maintaining material bounds throughout the optimization process.
The proposed method is evaluated against state-of-the-art Genetic Algorithm (GA) and inverse model baselines on synthetic and in vitro experimental datasets. Our approach demonstrates superior performance on synthetic data, achieving a Mean Absolute Error (MAE) of 0.06 compared to the baselines’ MAE of 0.18 and 0.54, respectively. On experimental specimens, our method outperforms the baseline in 5 out of 6 cases. Most importantly, our approach reduces calibration time to under three seconds, compared to up to 8 days per sample required by traditional calibration. Such efficiency paves the way for applying more complex FE models, enabling accurate patient-specific simulations and advancing spinal treatment planning.
@@
finite element model , intervertebral disc , surrogate model , neural network , calibration
@@
[label1]organization=Department of Diagnostic and Interventional Neuroradiology, Klinikum rechts der Isar, Technical University of Munich,city=Munich, country=Germany
@@
[label2]organization=Institute for AI in Medicine, School of Computation, Information and Technology, Technical University of Munich,city=Munich, country=Germany
@@ 1 Introduction
Finite element (FE) simulations are a well-established tool in orthopedics and spine research, providing significant value for diagnosing and planning treatments for conditions like scoliosis, fractures, degenerative disc disease, and osteoporosis [1]. Advances in patient-specific models improve surgical planning and enable personalized adjustments to treatment [1]. However, accurate numerical modeling of intervertebral discs (IVDs) encounters substantial challenges, such as structural complexity, patient-specific variability, and the necessity for selecting appropriate material definitions [2, 3].
Calibration or Parameter Estimation of an FE model ensures its accuracy and validity by finding input parameters such that the model output closely matches reported in vitro experimental measurements for particular specimens and geometries. This process is crucial because relying solely on material properties from the literature often fails to capture specific mechanical behaviors, particularly for varying geometries or defect conditions [4, 5]. Schlager et al. [6] reported high variability in material parameter definitions across IVD FE studies, stemming from differences in measurement methods used in in vitro experiments and distinct characteristics of the specimens [7]. This variability, together with the FE model’s complexity and the number of material parameters involved, also presents significant challenges for the calibration process.
@@
In a previous study [30, forthcoming], we demonstrated that using sensitivity analysis with an NN surrogate enhances the understanding of an FE model’s input features. This insight motivates us to guide the optimization with gradients derived from the surrogate, thereby increasing the efficiency of parameter estimation. Optimizing using NN gradients has been effective in other fields requiring input edits, such as ML robustness [31] and explainability [32]. In a different biomechanical domain, gradients from an NN surrogate guided a calibration of an FE model, incorporating an auxiliary optimization objective to implicitly promote feasible solutions [33].
The objectives of this work are two-fold. First, we implement and evaluate the performance of multiple ML models as surrogates for an established L4-L5 IVD FE model [14, 7]. To our knowledge, we are the first to develop an NN surrogate for an IVD FE model. Second, we introduce a novel, efficient, and effective calibration method guided by the NN surrogate using Projected Gradient Descent (PGD). Unlike other methods, our approach does not rely on external optimization algorithms [26, 21, 24, 27, 14] and explicitly ensures solution feasibility [33]. We compare our calibration method to a state-of-the-art GA [14] and a developed inverse model [28, 29] by evaluating it on a synthetic dataset and real-world experimental measurements, highlighting the practical applicability and limitations of the proposed approach. Our code is available at https://github.com/matanat/IVD-CalibNN/.
@@ 2 Materials and methods
Figure 1: Calibrating an L4-L5 intervertebral disc (IVD) finite element (FE) model to match in vitro Range-of-Motion (RoM) measurements. (1) Create a dataset by sampling material parameters within feasible bounds using Latin hypercube sampling (LHS) [34] and obtaining corresponding RoM values with FE simulations. (2) Train a neural network (NN) surrogate to minimize the Mean Absolute Error (MAE) between the predicted and simulated RoM. (3) Freeze the network weights and optimize the NN input parameters to match the predicted RoM to the experimental measurements. Projected Gradient Descent (PGD) ensures the optimized parameters remain within feasible bounds.
@@
We present a three-step method for calibrating an L4-L5 IVD FE model to match experimental Range-of-Motion (RoM) measurements (Fig. 1). First, we create a dataset using Latin hypercube sampling (LHS) [34] and generate corresponding FE simulation RoM outputs. Next, we train a surrogate NN model to approximate the FE model’s behavior, enabling efficient predictions. Finally, we calibrate the FE model to experimental measurements by optimizing the input parameters with PGD. During this calibration step, the weights of the pre-trained NN model are frozen, ensuring that only the inputs are adjusted and leveraging the surrogate’s learned representation.
@@
We begin this section by describing the FE model and the dataset used for training the surrogate models. Afterward, we detail the surrogate model architectures and the training process. Finally, we explain the calibration process, including the optimization method and the baseline comparisons to a GA and an inverse model.
2.1 FE model and dataset
Let $\mathbf{p}_{i}=[p_{1},\ldots,p_{d}]\in\mathbb{R}^{d}$ represent a set of material configurations, where $d=13$ denotes the dimensionality of the parameter space and $1\leq i\leq n$ indexes the individual configurations. The configurations were sampled with LHS, thus ensuring a comprehensive exploration of the parameter ranges by dividing them into equally probable intervals. Each input parameter $p_{j}$ was restricted to a feasible range [30, forthcoming], as detailed in B.1. Subsequently, we experimented with surrogates trained on configuration sets of various sizes.
The sampled configurations $\mathbf{p}$, each comprising 13 model parameters, were used to perform simulations with an established FE model of the human L4-L5 IVD [7, 14]. Besides $\mathbf{p}$, the simulation inputs a Load Case definition specifying the direction and a Moment magnitude of the applied mechanical loading. Specifically, load cases $c\in(1,2,3,4)$ were used, representing axial rotation, extension, flexion, and lateral bending, as detailed by Heuer et al. [9]. For each load case, five different moment magnitudes $m\in(1,2,3,4,5)$, measured in Nm, were applied. The output of each simulation was the rotational displacement of the IVD, expressed as the RoM in degrees $r_{c,m}\in\mathbb{R}_{>0}$. This RoM represents the rotational change around the axis corresponding to the applied load case, measured relative to the initial position.
@@ 2.1 FE model and dataset
Figure 2: Meshed geometry of the L4-L5 IVD and the strain energy density functions in the FE model [7, 14]. The figure shows the separation into nucleus pulposus (gray) and annulus fibrosus (blue), both with hyperelastic material definitions. The strain energy density functions for the nucleus ($\mathit{W_{n}}$, Mooney-Rivlin) and annulus ($\mathit{W_{a}}$, Holzapfel-Gasser-Ogden) are provided, with the latter specifically accounting for the anisotropic properties of the annulus. The calibration process involved 13 parameters, with five in the strain energy density functions (bold). The remaining parameters, used in functions not provided here, describe fiber dispersion, angle variations, and stiffness changes across the annulus. For a more detailed explanation, refer to our previous publications [7, 14].
The IVD geometry used in the FE model (Fig. 2) was manually reconstructed based on average dimensions reported in the literature [14]. The model included a clear separation between the nucleus pulposus (gray) and the annulus fibrosus (blue), both defined with hyperelastic material properties. The Mooney-Rivlin material model ($\mathit{W_{n}}$) described the material behavior of the nucleus pulposus, while the Holzapfel-Gasser-Ogden material model ($\mathit{W_{a}}$) [35] was used for the annulus fibrosus to account for its anisotropic properties [8]. These strain energy density functions, $\mathit{W_{n}}$ and $\mathit{W_{a}}$, describe the energy stored in a material per unit volume during deformation [36]. The annulus fibrosus was segmented into five symmetrical subregions, each containing five layers to reflect radial differences in the material properties of the IVD. The modeling process, combined with the specific material parameters from the strain energy density functions, resulted in 13 different parameters characterizing the behavior of the FE model. The model was meshed using quadratic hexahedral elements.
@@ 2.2 Surrogate models
Following the literature [15], we evaluated Linear Regression (LR), Support Vector Machine for Regression (SVR) [37], Random Forest (RF), Light Gradient-Boosting Machine (LightGBM) [38], Gaussian Process (GP), and NN surrogates. We used off-the-shelf implementations from scikit-learn [39] for LR, SVR, RF, and GP, and the official one for LightGBM [38]. The NN was implemented with PyTorch [40]. Hyperparameters for all models were searched with Optuna [41], with details provided in B.3. The models were trained and evaluated on the same machine mentioned before.
@@
We employed a simple feed-forward architecture for the NN consisting of five linear layers, each followed by a ReLU activation function. The final layer outputs the predicted RoM value and is not constrained to a specific range (e.g., (0,1)), allowing for extrapolation to unseen RoM values during calibration to some specimens. The loss function was the Mean Absolute Error (MAE) between the predicted and ground-truth RoM values.
@@ 2.3.1 Projected Gradient Descent with NN (PGD w/ NN)
The pre-trained NN surrogate guides the search for input parameters $\mathbf{x}_{\text{cal}}\in\mathbb{R}^{d}$ to achieve desired RoM values $\mathbf{y}\in\mathbb{R}^{k}$ for all load cases and moments specified by an experimental measurement, i.e., $k=|LoadCases|\times|Moments|$. We employed Projected Gradient Descent (PGD) [43, p. 263] to optimize $\mathbf{x}_{\text{cal}}$ so that the network output matches $\mathbf{y}$. During this process, the NN weights were frozen and not changed, ensuring that the model’s learned representation remained consistent and that only the input parameters were adjusted. This process differs from the one used during the training phase of the NN, where the weights were updated to learn the mapping from inputs to outputs.
@@ 2.3.2 Calibration baselines
We employed a GA [7, 14] as a first baseline, substituting the FE model in the original works with the NN surrogate. We opted for this approach because running the calibration without the surrogate would be infeasible due to the excessive running time. We used the settings from the original works [7, 14] while restricting the sampled parameters bounds as described previously and increasing the $\overline{\mathcal{R}^{2}}$ stopping criteria to 1. We increased this criteria for fairer comparison with the developed method and since using the surrogate circumvents previous computational restrictions. Additional details are provided in B.2.
@@ 3 Results and discussion
This section compares the performance of the different surrogate models used to approximate FE model simulations. Then, we discuss the results of the calibration methods by evaluating them on both synthetic and experimental data. Finally, we conduct ablation studies to understand the impact of various components of our approach. Additional results are provided in C.
@@ 3.1 Surrogate models
To enhance the efficiency and accuracy of FE calibration, we explore using surrogate models as a faster and more reliable alternative to direct simulations. This section assesses the performance of various surrogate models, focusing on the effects of training dataset size and model architecture. Additionally, it investigates the models’ ability to interpolate and extrapolate RoM values to unseen moments, which is essential for calibrating to experimental data.
3.1.1 Training set size and model architecture
The size of the dataset used for training surrogate models is critical due to the time-consuming nature of FEM simulations. Table 1 explores the number of samples required for accurate model predictions, comparing datasets of 128, 512, and 1024 samples, all generated using LHS. The models were evaluated on a test set of 64 randomly sampled configurations. All ground-truth RoM values were obtained by running FE simulations, and all experiments were conducted with 4-fold cross-validation.

### v2

# Neural Network Surrogate and Projected Gradient Descent for Fast and Reliable Finite Element Model Calibration: a Case Study on an Intervertebral Disc

@@ Abstract
Accurate calibration of finite element (FE) models is essential across various biomechanical applications, including human intervertebral discs (IVDs), to ensure their reliability and use in diagnosing and planning treatments. However, traditional calibration methods are computationally intensive, requiring iterative, derivative-free optimization algorithms that often take days to converge.
This study addresses these challenges by introducing a novel, efficient, and effective calibration method demonstrated on a human L4-L5 IVD FE model as a case study using a neural network (NN) surrogate. The NN surrogate predicts simulation outcomes with high accuracy, outperforming other machine learning models, and significantly reduces the computational cost associated with traditional FE simulations. Next, a Projected Gradient Descent (PGD) approach guided by gradients of the NN surrogate is proposed to efficiently calibrate FE models. Our method explicitly enforces feasibility with a projection step, thus maintaining material bounds throughout the optimization process.
The proposed method is evaluated against state-of-the-art Genetic Algorithm (GA) and inverse model baselines on synthetic and in vitro experimental datasets. Our approach demonstrates superior performance on synthetic data, achieving a Mean Absolute Error (MAE) of 0.06 compared to the baselines’ MAE of 0.18 and 0.54, respectively. On experimental specimens, our method outperforms the baseline in 5 out of 6 cases. While our approach requires initial dataset generation and surrogate training, these steps are performed only once, and the actual calibration takes under three seconds. In contrast, traditional calibration time scales linearly with the number of specimens, taking up to 8 days in the worst-case. Such efficiency paves the way for applying more complex FE models, potentially extending beyond IVDs, and enabling accurate patient-specific simulations.
@@
finite element model , calibration , surrogate , neural network , intervertebral disc
@@
[label1]organization=Institute for Neuroradiology, TUM University Hospital, School of Medicine and Health, Technical University of Munich (TUM),city=Munich, country=Germany
@@
[label2]organization=Chair for AI in Healthcare and Medicine, Technical University of Munich (TUM) and TUM University Hospital,city=Munich, country=Germany
@@
[…]
[label7]organization=Department of Computing, Imperial College London, city=London, country=United Kingdom
@@ 1 Introduction
Finite element (FE) simulations are a well-established tool across various fields in biomechanics and are integral to applications in orthopedics, tissue modeling, and spinal research [1, 2, 3, 4]. However, accurate numerical modeling encounters substantial challenges, such as structural complexity, diverse material behaviors, and patient-specific variability [5, 6].
One of the most demanding aspects of FE modeling is Calibration or Parameter Estimation, which ensures model accuracy by adjusting input material parameters until the model output closely matches in vitro experimental or in vivo patient-specific measurements [7]. This process is crucial because relying solely on literature-based material properties often fails to capture the specific mechanical behaviors of tissues, such as intervertebral discs (IVDs), under varying geometries or defect conditions [8, 9]. In IVD modeling, for instance, calibration is necessary to overcome the high variability and uncertainty in material parameter definitions across studies, which arises from differences in measurement methods and specimen characteristics [10, 11, 12]. This challenge is not limited to IVD models, as uncertainties in material parameters and model complexity affect the accuracy of FE simulations in various biomechanical applications [13, 14]. Therefore, calibration is essential to make FE models more reliable for clinical application.
@@
Gradients derived from NNs have been effectively employed in optimization problems across various domains, such as improving model robustness [36] and explaining its predictions [37]. Inspired by these approaches, we leverage gradients from the surrogate model to guide the calibration process, thereby improving the efficiency of parameter estimation. In biomechanics, NN surrogate gradients have been used to calibrate a left ventricle FE model, incorporating an auxiliary optimization objective that implicitly supports feasibility [38]. In contrast, our approach explicitly enforces feasibility, setting it apart from this prior work.
In this study, we focus on the L4-L5 IVD as a case study due to its clinical relevance, nonlinear behavior, significant motion, and exposure to high loads [39, 40, 41, 42]. This IVD segment is often affected by degenerative diseases, making it a common target for surgical interventions and biomechanical research.
The objectives of this work are two-fold. First, we implement and evaluate the performance of multiple ML models as surrogates for an established L4-L5 IVD FE model [20, 10]. To our knowledge, we are the first to develop an NN surrogate for an IVD FE model. Second, we introduce a novel, efficient, and effective calibration method guided by the NN surrogate using Projected Gradient Descent (PGD). Unlike other methods that rely on independent optimization frameworks, such as mesh-based solvers or GAs, treating the surrogate as a black box [32, 27, 30, 33, 20], our approach directly leverages gradients from the surrogate, enabling efficient calibration while explicitly ensuring solution feasibility [38]. We compare our calibration method to a state-of-the-art GA [20] and a developed inverse model [34, 35] by evaluating it on a synthetic dataset and real-world experimental measurements, highlighting the practical applicability and limitations of the proposed approach. Our code is available at https://github.com/matanat/IVD-CalibNN/.
@@ 2 Materials and methods
Figure 1: Calibrating an L4-L5 intervertebral disc (IVD) finite element (FE) model to match in vitro measurements. (1) Create a dataset by sampling material parameters within feasible bounds using Latin hypercube sampling (LHS) [43] and obtaining corresponding range of motion (RoM) values with FE simulations. (2) Train a neural network (NN) surrogate to minimize the Mean Absolute Error (MAE) between the predicted and simulated RoM. (3) Freeze the network weights and optimize the NN input parameters to match the predicted RoM to the measurements. Projected Gradient Descent (PGD) ensures the optimized parameters remain within feasible bounds. Finally, (4) validate the calibrated material parameters by comparing simulated intradiscal pressure (IDP) to experimental measurements.
@@
We present a method for calibrating an FE model to match experimental measurements and demonstrate its applicability on an L4-L5 IVD FE (Fig. 1). First, we create a dataset using Latin hypercube sampling (LHS) [43] and generate corresponding FE simulation range of motion (RoM) outputs. Next, we train a surrogate NN model to approximate the FE model’s behavior, enabling efficient predictions. We calibrate the FE model to experimental measurements by optimizing the input parameters with PGD. During this calibration step, the weights of the pre-trained NN model are frozen, ensuring that only the inputs are adjusted and leveraging the surrogate’s learned representation. Finally, we validate the calibrated parameters by comparing their respective simulated intradiscal pressure (IDP) to experimental measurements.
@@
We begin this section by describing the FE model and the dataset used for training the surrogate models. Afterward, we detail the surrogate model architectures and the training process. Finally, we explain the calibration process, including the optimization method, the baseline comparisons to a GA and an inverse model and the validation process.
2.1 FE model
We used an established FE model of the human L4-L5 IVD [10, 20]. The IVD geometry used in the FE model (Fig. 2) was manually reconstructed based on average dimensions reported in the literature. The model separates the nucleus pulposus (gray) and the annulus fibrosus (blue), both characterized by hyperelastic material properties, with the nucleus pulposus comprising 44% of the total disc volume. The annulus fibrosus was segmented into five symmetrical subregions, each containing five layers to reflect radial differences in the material properties of the IVD. The Mooney-Rivlin material model ($\mathit{W_{n}}$) described the material behavior of the nucleus pulposus, while the Holzapfel-Gasser-Ogden material model ($\mathit{W_{a}}$) [44] was used for the annulus fibrosus to account for its anisotropic properties resulting from its collagen fibers [7]. These strain energy density functions, $\mathit{W_{n}}$ and $\mathit{W_{a}}$, describe the energy stored in a material per unit volume during deformation [45]. The IVD was simulated without adjacent vertebrae, with a reference point 10 mm above the center of the disc’s cross-sectional surface for load application [10]. A coupling constraint connected this point to the upper surface of the IVD, while the lower surface was fixed via a coupling constraint to an anchored reference point, replicating the fixed caudal vertebra in in vitro experiments. The model was meshed using quadratic hexahedral elements.
The modeling process, combined with the specific material parameters from the strain energy density functions, resulted in 13 input parameters characterizing the behavior of the FE model.
@@ 2.1 FE model
Figure 2: Meshed geometry of the L4-L5 IVD and the strain energy density functions in the FE model [10, 20]. The figure shows the separation into nucleus pulposus (gray) and annulus fibrosus (blue), both with hyperelastic material definitions. The strain energy density functions for the nucleus ($\mathit{W_{n}}$, Mooney-Rivlin) and annulus ($\mathit{W_{a}}$, Holzapfel-Gasser-Ogden) are provided, with the latter specifically accounting for the anisotropic properties resulting from the collagen fibers in the annulus. The calibration process involved 13 parameters, with five in the strain energy density functions (bold). The remaining parameters, used in functions not provided here, describe fiber dispersion, angle variations, and stiffness changes across the annulus. For a more detailed explanation, refer to our previous publications [10, 20].
2.2 Dataset creation
As a preprocessing step, we considered a global sensitivity analysis of the same FE model, performed in a parallel work [46]. This analysis revealed that some input parameters had significant influence while others had minimal effect. However, the influence of most parameters varied across load scenarios. To avoid missing important interactions or non-linear effects, we decided to include all 13 parameters in the surrogate model training setting.
Let $\mathbf{p}_{i}=[p_{1},\ldots,p_{d}]\in\mathbb{R}^{d}$ represent a set of material configurations, where $d=13$ denotes the dimensionality of the parameter space and $1\leq i\leq n$ indexes the individual configurations. The configurations were sampled with LHS, thus ensuring a comprehensive exploration of the parameter ranges by dividing them into equally probable intervals. Each input parameter $p_{j}$ was restricted to a feasible range as detailed in B.1. Subsequently, we experimented with surrogates trained on configuration sets of various sizes.
The sampled configurations $\mathbf{p}$, each comprising d parameters, were used to perform simulations with the FE model. Besides $\mathbf{p}$, the simulation inputs a Load Case definition specifying the direction and a Moment magnitude of the applied mechanical loading. Specifically, load cases $c\in(1,2,3,4)$ were used, representing axial rotation, extension, flexion, and lateral bending, as detailed by Heuer et al. [15]. For each load case, five different moment magnitudes $m\in(1,2,3,4,5)$, measured in Nm, were applied. The output of each simulation was the rotational displacement of the IVD, expressed as the RoM in degrees $r_{c,m}\in\mathbb{R}_{>0}$. This RoM represents the rotational change around the axis corresponding to the applied load case, measured relative to the initial position.
@@ 2.3 Surrogate models
Following the literature [21], we evaluated Linear Regression (LR), Polynomial Regression (PR), Support Vector Machine for Regression (SVR) [47], Random Forest (RF), Light Gradient-Boosting Machine (LightGBM) [48], Gaussian Process (GP), and NN surrogates. We used off-the-shelf implementations from scikit-learn [49] for LR, PR, SVR, RF, and GP, and the official one for LightGBM [48]. The NN was implemented with PyTorch [50]. Hyperparameters for all models were searched with Optuna [51], with details provided in B.3. The models were trained and evaluated on the same machine mentioned before.
@@
We employed a simple feed-forward architecture for the NN consisting of five linear layers, each followed by a ReLU activation function. The final layer outputs the predicted RoM value and is not constrained to a specific range (e.g., (0,1)), allowing for later extrapolation to unseen RoM values during calibration for some specimens. The loss function was the Mean Absolute Error (MAE) between the predicted and ground-truth RoM values.
@@ 2.4.1 Projected Gradient Descent with NN (PGD w/ NN)
The pre-trained NN surrogate guides the search for input parameters $\mathbf{x}_{\text{cal}}\in\mathbb{R}^{d}$ to achieve desired RoM values $\mathbf{y}\in\mathbb{R}^{k}$ for all load cases and moments specified by an experimental measurement, i.e., $k=|LoadCases|\times|Moments|$. We employed Projected Gradient Descent (PGD) [53, p. 263] to optimize $\mathbf{x}_{\text{cal}}$ so that the network output matches $\mathbf{y}$. During this process, the NN weights were frozen and not changed, ensuring that the model’s learned representation remained consistent and that only the input parameters were adjusted. This process differs from the one used during the training phase of the NN, where the weights were updated to learn the mapping from inputs to outputs.
@@ 2.4.2 Calibration baselines
We employed a GA [10, 20] as a first baseline, substituting the FE model in the original works with the NN surrogate. We opted for this approach because running the calibration without the surrogate would be infeasible due to the excessive running time. We used the settings from the original works [10, 20] while restricting the sampled parameters bounds as described previously and increasing the $\overline{\mathcal{R}^{2}}$ stopping criterion to 1. We increased this criterion for fairer comparison with the developed method and since using the surrogate circumvents previous computational restrictions. Additional details are provided in B.2.
@@
2.5 Validation
Validation is performed to ensure that the computational model reflects known physiological behavior [54]. In this study, we compared simulated IDP, which was not part of the calibration process, to ranges reported in the literature [55]. In the simulation, IDP was measured at the center of the nucleus pulposus in megapascal (MPa), with observed values ranging between 0 and 2.5 MPa. In contrast to the load cases used for training the surrogate, validation was performed using axial compression loading with an applied compressive force ranging from 0 to 2000 Newtons (N).
@@ 3 Results and discussion
This section compares the performance of the different surrogate models used to approximate FE model simulations. We discuss the results of the calibration methods by evaluating them on both synthetic and experimental data. Then, we present the results of the validation process. Finally, we conduct ablation studies to understand the impact of various components of our approach. Detailed results are provided in C.
@@ 3.1 Surrogate models
To enhance the efficiency and accuracy of FE calibration, we explore using surrogate models as a faster and more reliable alternative to direct simulations. This section assesses the performance of various surrogate models, focusing on the effects of model architecture and training dataset size. Additionally, it investigates the models’ ability to interpolate and extrapolate RoM values to unseen moments, which is essential for calibrating to experimental data. The following experiments were designed to identify the best-performing surrogate model with rigorous evaluation on large test sets, where the predicted RoM values are compared to ones obtained from FE simulations. In the subsequent sections, the best-performing surrogate was evaluated on its ability to calibrate material parameters with experimental data, where only a limited number of samples are available.
3.1.1 Model architecture
To assess the performance of various surrogate model architectures, we trained the models on a dataset of 1024 LHS samples and evaluated them on a test set of 64 randomly sampled configurations. All experiments were conducted with 4-fold cross-validation.
In the last two columns of Table 1, models commonly used in Response Surface methods [57, 23], such as LR and PR, result in poor performance. In particular, the PR models show improvement with increasing polynomial degrees of 4 and 5; however, they remain prone to overfitting, as indicated by their high MAE (0.49 and 0.53, respectively). More advanced methods, such as GP and LightGBM, show stronger performance (MAE of 0.33 and 0.28). Nevertheless, the NN outperforms all others, with an MAE of 0.10 and $\overline{\mathcal{R}^{2}}$ of 0.99, highlighting its superior ability to handle the non-linear interactions in the FE simulations [10]. As surrogate accuracy is crucial for the calibration task, the NN was chosen for the subsequent experiments.